Find the derivatives of the following functions.



• y=(4x+3)/ln[2x^2+3x-5]

1
Expert's answer
2022-01-02T18:19:16-0500

"y=\frac{4x+3}{\ln(2x^2+3x-5)},"

"y'=\frac{4\ln(2x^2+3x-5)-\frac{(4x+3)^2}{2x^2+3x-5}}{\ln^2(2x^2+3x-5)}."